OpenStudy (anonymous):

find the integral: x^3+5/x dx

OpenStudy (anonymous):

\[\int \frac{x^3+5}{x}dx\]or \[\int( x^3+\frac{5}{x})dx\]?

OpenStudy (anonymous):